OpenStudy (anonymous):

The length of a rectangle deck is 8 meters more than twice the width. The perimeter of the deck is 46m. Find the length. HELP~

OpenStudy (anonymous):

L=8+2W 2L+2W=46 Solve by substitution: 2(8+2W)+2W=46 16+4W+2W=46 16+6W=46 6W=30 W=5 Then plug in and solve for L: 8+2(5) 8+10 L=18

OpenStudy (anonymous):

omg thank thank thank you !

OpenStudy (anonymous):

you're welcome :)